### v1

# Collision Energy Dependence of Particle Ratios and Freeze-out Parameters in Ultra Relativistic Nucleus Nucleus Collisions

Abstract
This work investigates how the theoretically calculated particle ratios and freeze-out parameters for a multi-component hot and dense non-ideal hadron resonance gas system, formed during the pre freeze-out stage in ultra-relativistic nucleus-nucleus collisions, are affected by van der Waals type interactions, incorporating both repulsive and attractive interactions among hadrons. The baryons (antibaryons) are treated as hard incompressible objects. Using this theoretical approach, we analyze experimental particle ratios involving like-mass as well as unlike-mass hadrons. We also try to understand the existence of horn-like structures seen in the ratios of strange particles to pion as a function of the collision energy. Using information about the dependence of experimental particle ratios on $\sqrt{s_{NN}}$ over a wide range of collision energies, we have extracted several model freeze-out parameters and examined their dependence on $\sqrt{s_{NN}}$ The correlation between the $\bar{p}/p$ and $K^{-}/K^{+}$ ratios is also examined.
22footnotetext: Corresponding Author E-mail: suddin@jmi.ac.in11footnotetext: E-mail: reshiiqbal24@gmail.com
I INTRODUCTION
Understanding the quark-hadron phase transition diagram and the nuclear matter under extreme conditions also requires investigating the thermodynamics of the hot and dense matter created in ultra-relativistic nucleus-nucleus collisions (URNNC). The URNNC are a key tool for experimentally studying the high energy-density and high-temperature states of matter, such as the Quark-Gluon Plasma (QGP). These collisions provide insights into collective effects, strange hadron yields and the possibility of a potential phase transition from a hadron resonance gas (HRG) phase, where quarks and gluons are confined to a phase where they are deconfined or chirally symmetric phase. The URNNC programs at the Relativistic Heavy Ion Collider (RHIC) at BNL and the Large Hadron Collider (LHC) at CERN aim to study the properties of strongly interacting matter under extreme conditions. They also offer a unique opportunity to study the chemical equilibration of hadrons in the final state of collisions Muller1995 ; Adcox2005 ; Singh1992 ; Tannenbaum2006 ; Cleymans2011 .
Understanding the behavior of matter a few microseconds after the Big Bang in the early universe and within the core of the neutron stars has gained significant attention. In URNNC the exploration of strongly interacting matter at extreme temperatures and densities is a focal point for both experimental and theoretical investigations. At the laboratory level experimental endeavors such as those conducted at the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC) play a crucial role in unraveling the mysteries of hot and dense hadronic matter. The study of URNNC has shown that statistical-thermal models can quite successfully reproduce the essential particle production features and their relative multiplicities in such collisions. The statistical hadronization model (SHM) also reproduces transverse momentum spectra of different particle species, making it a useful tool for analyzing soft hadron production, particularly strangeness production which has been proposed as a signature of QGP formation Metag1993 ; Cleymans1986O ; Letessier1993 ; Becattini1998 ; Munzinger1999 ; Yen1998 ; Parra2019S .
Theoretical models, particularly the HRG model, have been instrumental in providing insights into the complex dynamics of strongly interacting particles. The Ideal Hadron Resonance Gas (IHRG) model, an initial formulation, treats mesons and baryons as point-like entities, assuming no interaction among them. However, deviations from an uncorrelated gas of hadrons are observed in lattice Quantum Chromodynamics (QCD) simulations Becattini2000 ; Munzinger2001 ; Florkowski2001 ; Becattini1995 ; Cleymans1999 ; Munzinger1994 . The ideal excluded volume type HRG models play a crucial role in understanding the quark-hadron phase transition and the particle production in URNNC Andronic2006 . In a Monte Carlo type approach with excluded-volume effect taken into account in the HRG, calculation of arbitrary moments of the event-by-event hadron yields have been done. However an analysis of different types of particle ratios is still lacking Bazavov2012 . It nevertheless highlights the systematic excess of experimental pion abundances compared to ideal gas results, which attempted by assigning the pions a small hard-core volume Borsanyi2014 . These can lead to further review the various thermal and statistical descriptions of particle production in URNNC, including the excluded-volume effect in the HRG model.
Standard thermal model have generally no interactions among the hadrons after a stage called freeze-out. However, with no interaction taken into account during the evolution of the HRG till its final breakup stage, one cannot understand the hadron production in a realistic manner in URNNC. Hence, we need thermal models which incorporate interactions to study the properties of a hot and dense HRG including the relative hadronic yields in such a system Vovchenko2018M . To address these challenges within the framework of a HRG model we can incorporate the attractive as well as the short-range repulsive interactions in van der Waals (VDW) type approach. Its ability to describe the nuclear liquid-gas transition at low temperatures and high baryon densities can set the stage for further useful exploration of the thermodynamic properties of HRG Yen1997 ; Tiwari2013 ; Mishra2008 ; Singh1991 ; Bashir2019P ; Das2016C .
For hot and dense hadron gas, several approaches have been proposed to formulate a precise equation of state (EOS), often using phenomenological models due to the lack of lattice QCD results at finite baryon density $(n_{B})$. Thermal models, assuming thermo-chemical equilibrium, are effective in characterizing HRG, with the IHRG model being a basic version that treats hadrons and resonances as non-interacting point-like particles at freeze-out Welke1990 . It has been established that the short-range repulsion must be included to accurately describe HRG behaviour, as the IHG model’s phase transition calculations reveal inconsistencies at high temperatures and baryon densities Cleymans1986O ; Cleymans1986Q . The hadronic hard-core repulsive interactions become significant in a dense HRG and hence in a macroscopic approach this can be taken into account by assigning hadrons a geometrical hard-core volume, resulting in a van der Waals excluded-volume effect that reduces particle densities in the system. Recent phenomenological models incorporating this effect Singh1991 ; Cleymans1987I ; Kouno1989 ; Rischke1991 ; Anchishkin1992 ; Mir2023 ; Uddin2012 ; Parra2019 ; Mir2024 ; Bashir2018I ; Parra2022M often lack a well-defined partition function or thermodynamic potential ($\Omega$), leading to thermodynamical inconsistencies in the calculations of particle densities, i.e. $n\neq\partial\Omega/\partial\mu$, where $\mu$ is the chemical potential of the particle and. It also breaks causality in dense HRG, making them less appealing despite their some empirical success.
Some earlier studies have also highlighted the importance of studying particle ratios as a function of the centre of mass energy in the URNNC. Hanafy (2021) Hanafy2021 found that a statistically corrected HRG model can reasonably describe experimental results across different energies. Tawfik (2012) Tawfik2012P also emphasized the significance of these ratios, particularly in the context of dynamical fluctuations in URNNC. Earlier Bjorken (1974) Bjorken1974 had also underscored the relevance of these ratios in understanding particle production in energetic collisions Blume2011 ; Tiwari2011 ; Tawfik2017 ; Stiles2006 ; Mir2024 ; Andronic2005 ; Bashir2015 ; Chatterjee2015 ; Mishra2008 ; Tiwari2013 .
The interesting horn-like structure in the kaon to pion ratio, as observed in URNNC, has been attributed to the release of a large number of colour degrees of freedom Nayak2005 ; Nayak2010 ; Nayak2011 . This phenomenon has been further explored in a hadronic kinetic model, which successfully reproduces the sharp peak in the $K^{+}/\pi^{+}$ ratio Tomik2006 . These findings collectively suggest that the horn in the kaon to pion ratio is a result of the complex interplay of colour degrees of freedom and collision processes. However, in the present study we have attempted to investigate in detail this characteristic horn structure in the kaon-to-pion ratio as well as other strange hadron to pion ratios in URNNC. Within the framework of the model used in the present study we have found strong evidence of strangeness imbalance factor, $\xi_{s}$. This helps us in understanding the production of strange particles relative to pions in the system produced in URNNC Bir1982 ; Grebieszkow2009 ; Bhattacharyya2017 ; Tawfik2021H ; Cleymans2017T ; Oeschler2017 ; Cleymans2004T .
The present work is organized as follows. In Section II we have provided some basic idea of the model formulation. In Section III the results of our calculations and their comparison with the experimental data is presented along with the discussion. Finally in Section IV we have summarized our results.
II STATISTICAL THERMAL MODEL
A simple and popular model for describing the pressure function for a system of particles in equilibrium having both repulsive and attractive interactions is the one described by the van der Waals (VDW) equation of state. In canonical ensemble representation, it can be written as Samanta2017C ; Greiner1965 ; Landau1980 :
$\left[P+\left(\frac{N}{V}\right)^{2}a\right]\left(V-b\right)=NT$ (1)
where the attractive and repulsive interactions between the particles are described, respectively, by the positive VDW parameters $a$ and $b$. The pressure, volume, temperature, and the number of particles having an “effective finite-size” due to their hard-core repulsion among them in the system are represented by the letters $P$, $V$, $T$, and $N$, respectively. The above can be rewritten in a simple and transparent form as:
$p\left(T,n\right)=\frac{NT}{V-bN}-a\frac{N^{2}}{V^{2}}=\frac{nT}{1-bn}-an^{2}$ (2)
The symbol $n$ represents the finite-size particle number density. The first term includes the correction factor due to hard-core repulsion by assigning a hard-core radius to the particles (baryons and antibaryons in our case). Considering particles as rigid spheres, we have $b=16\pi r^{3}/3$, where $r$ is the particle’s hard-core radius. The second term accounts for the attractive interactions between particles through the parameter $a$. For $a=0$, the equation reduces to the EV-HRG equation, containing only repulsive interactions, while for both $a=0$ and $b=0$ it reduces to the ideal HRG case Greiner1965 ; Tiwari2011 ; Bugaev2000 ; Sahoo2023 ; Yen1997 .
In a statistical approach, it is necessary to adopt the grand canonical ensemble (GCE) formulation for the system Samanta2017C ; Vovchenko2018Q ; Vovchenko2017M ; Pradhan2022 ; Vovchenko2015P ; Sarkar2018 ; Vovchenko2015S ; Vovchenko2015V . The preference for the Grand Canonical Ensemble (GCE) over the Canonical Ensemble (CE) stems from its technical simplicity in describing the properties of the systems due to its non-conserved (fluctuating) particle numbers and energy. Hence, the GCE is most suited for the HRG system formed in the URNNC essentially due to the particle production/annihilation reactions taking place inside it. The CE formulation thus lacks a comprehensive thermodynamic description for such systems due to its dependence on parameters like volume ($V$), temperature ($T$), energy ($E$), and particle number ($N$). In the GCE formulation, we can calculate mean particle number density ($n$), energy density ($\varepsilon$), entropy density ($s$), etc., which are essential for understanding the system’s behaviour. Here we will apply the following statistical approach, which was earlier performed for the EV model, i.e., for $a=0$ in Eq. (2), Rischke1991 ; Gorenstein1981 ; Parra2019 ; Parra2023 ; Parra2019T ; Din2024 .
II.1 VDW EQUATION OF STATE THROUGH GRAND CANONICAL ENSEMBLE:
We discuss the grand canonical ensemble (GCE) formulation by considering both attractive and repulsive interactions. The repulsive force is assumed to exist between pairs of two baryons and pairs of two antibaryons, while it is purely attractive between a baryon-antibaryon pair as the annihilation processes predominantly govern short-range attractive interactions between baryon-antibaryon pairs. The hard-core repulsive interaction among bosons (mesons) has been neglected Vovchenko2016V ; Vovchenko2015V ; Samanta2017C ; Andronic2012 .
The grand canonical partition function ($Z_{GC}$) for a system can be written as:
$Z_{GC}\left(T,\mu,V\right)=\sum_{N=0}^{\infty}e^{\mu N/T}Z_{C}\left(T,N,V\right)$ (3)
where $Z_{C}\left(T,N,V\right)$ stands for the canonical partition function applicable to a system with $N$ number of particles. $T$ signifies the temperature and $V$ represents the total physical volume of the system. The phenomenological grand partition function, incorporating both attractive and repulsive interactions, can be formulated as follows:
$Z_{GC}^{int}\left(T,\mu,V\right)=\sum_{N=0}^{\infty}e^{\mu N/T}Z_{C}\left(T,N,V-bN\right)e^{-\bar{U}/T}$ (4)
Here, $Z_{C}\left(T,N,V-bN\right)$ is the canonical partition function accounting for the excluded volume effect due to hard-core repulsion. The term $bN$ represents the total excluded volume resulting from the hard-core repulsive interaction among baryons (or among antibaryons) Rischke1991 ; Uddin1994 ; Kouno1989 . The attractive interaction is incorporated by introducing the factor $e^{-\bar{U}/T}$ in the grand partition function, where $\bar{U}$ represents the average attractive interaction energy which can be expressed as
$\bar{U}=\frac{1}{2}\sum_{i,j}V_{att}\left(\vec{r_{i}}-\vec{r_{j}}\right)$ (5)
The summation over the indices $i,j$ runs over all particles ($N$) in the system. The $V_{att}\left(\vec{r_{i}}-\vec{r_{j}}\right)$ represents the mean attractive interaction energy between any two particle pairs. Interactions involving three or more particles are disregarded due to their rarity. Assuming a uniform finite-size particle number density, $n=\frac{N}{V}$, the total interaction energy can be determined by summing over all particle pairs. Therefore, for large $n$, we can express it in the integral form as followsLandau1980 ; Armiento2005 ; Ramshaw1980 :
$\bar{U}=\frac{1}{2}\int d^{3}\vec{r_{1}}d^{3}\vec{r_{2}}\,n\left(\vec{r_{1}}\right)n\left(\vec{r_{2}}\right)V_{att}\left(\vec{r_{2}}-\vec{r_{1}}\right)$ (6)
[…]
$Z_{GC}^{int}\left(T,\mu,V\right)=\sum_{N=0}^{\infty}e^{\mu N/T}Z_{C}\left(T,N,V-bN\right)e^{-n^{2}Va/T}$ (7)
Defining the grand partition function involving only the hard-core repulsive interaction as:
$Z_{GC}^{excl}\left(T,\mu,V\right)=\sum_{N=0}^{\infty}e^{\mu N/T}Z_{C}\left(T,N,V-bN\right)$ (8)
we can write:
$Z_{GC}^{int}\left(T,\mu,V\right)=Z_{GC}^{excl}\left(T,\mu,V\right)e^{-\frac{n^{2}Va}{T}}$ (9)
In order to define the system’s pressure, we adopt a standard procedure wherein we take the Laplace transformation of $Z_{GC}^{int}\left(T,\mu,V\right)$ Rischke1991 ; Parra2019 ; Parra2023 :
[…]
In order to overcome the extreme right-hand singularity, the above integral must be convergent in the limit $V\rightarrow\infty$. This provides:
$\zeta=\lim_{V\rightarrow\infty}\left[\frac{\ln Z_{GC}^{int}\left(T,\mu,V\right)}{V}\right]$ (11)
or
$\zeta=\frac{1}{T}\lim_{V\rightarrow\infty}\left[\frac{T\ln Z_{GC}^{int}\left(T,\mu,V\right)}{V}\right]$ (12)
This gives:
$\zeta=\frac{1}{T}p^{int}\left(T,\mu,V\right)$ (13)
We can again write the Laplace transform using Eq. 9 as:
[…]
Rewriting:
[…]
and
$\zeta=\frac{1}{T}\lim_{V\rightarrow\infty}\left[\frac{T\ln Z_{GC}^{excl}\left(T,\mu,V\right)}{V}-n^{2}a\right]$ (16)
This leads to:
$p^{int}\left(T,\mu,V\right)=p^{excl}\left(T,\mu,V\right)-an^{2}$ (17)
When a=0, the $p^{excl}\left(T,\mu,V\right)$ for the case of GCE turns out to be: Gorenstein1999V ; Rischke1991 ; Zeeb2002E
$p^{excl}\left(T,\mu,V\right)=\frac{nT}{1-bn}$ (18)
where $n$ in this case is the GCE averaged particle number density. This gives for Eq. 17 the following:
$p^{int}\left(T,\mu,V\right)=\frac{nT}{1-bn}-an^{2}$ (19)
In principle, the particle number density can be obtained in a thermodynamically consistent manner as $n=\left.\frac{\partial p^{int}}{\partial\mu}\right|_{T}$ Parra2019 ; Parra2023 :
$n=\left(\frac{T}{\left(1-bn\right)^{2}}-2an\right)\left.\frac{\partial n}{\partial\mu}\right|_{T}$ (20)
After further calculations, one can obtain an expression for the chemical potential, $\mu$, as:
$\mu=T\ln\left(\frac{n}{1-bn}\right)+\frac{T}{1-bn}-2an+C$ (21)
For $a=0,b=0$ we should recover the ideal case, i.e., $n(T,\mu)=n_{\mathrm{id}}(T,\mu)$ where we can write (using Boltzmann approximation) $n_{\mathrm{id}}\left(T,\mu\right)=\phi\left(T,m\right)\exp\left(\frac{\mu}{T}\right)$ for the point-like particle number density. The integration constant $C$ comes out to be Gorenstein2022 ; Vovchenko2015P ; Parra2023 ; Kostyuk2000 .
$C=T\ln\left(\frac{n_{\mathrm{id}}}{\phi}\right)-T\ln n_{\mathrm{id}}-T$ (22)
Thus we can write:
$\mu=T\ln\left(\frac{n}{(1-bn)\phi}\right)+\frac{Tnb}{1-bn}-2an$ (23)
Defining:
$\mu^{\ast}=T\ln\left[\frac{n}{(1-bn)\phi}\right]$ (24)
and again writing the quantity $\phi$ in terms of $\mu^{\ast}$ as:
$\phi=n^{\mathrm{id}}\left(T,\mu^{\ast}\right)e^{-\frac{\mu^{\ast}}{T}}$ (25)
we get:
$\mu^{\ast}=\mu-\frac{bnT}{1-bn}+2an$ (26)
Using Eq. 14 we can finally write the above result as:
$\mu^{\ast}=\mu-bp^{excl}\left(T,\mu,V\right)+2an$ (27)
Combining Eq. 18 and Eq. 19 we will get:
$n(T,\mu)=(1-bn)n^{\mathrm{id}}\left(T,\mu^{\ast}\right)$ (28)
which finally yields for the number density of the particle $n\left(\mu,T\right)=n^{int}(T,\mu)$ as:
$n\left(\mu,T\right)=\frac{n^{\mathrm{id}}\left(T,\mu^{\ast}\right)}{1+bn^{\mathrm{id}}\left(T,\mu^{\ast}\right)}$ (29)
In view of the result in Eq. 19, the above Eq. 25 takes the form of a transcendental equation. For a multi-component system, it is possible to generalize the above result for a given $j$th particle species as Zeeb2002E :
$n_{j}\left(\mu_{j},T\right)=\frac{{n^{\mathrm{id}}}_{j}\left(T,\mu_{j}^{\ast}\right)}{1+\sum_{i}{b{n^{\mathrm{id}}}_{i}\left(T,\mu_{j}^{\ast}\right)}}$ (30)
The summation over the index $i$ is for all species having hard-core repulsive interaction, including the $j$th specie. Further details can be found in Parra2019 . The above results can be effectively utilized to calculate relative particle yields in the system.
The resonances that decay into specie $j$ also contribute to its finally measured yield in the actual experiments. Therefore, the contributions from all heavier hadrons (say $k^{th}$ up to the mass 2 GeV) that decay into hadron $j$ with their known branching fraction $\Gamma_{k\rightarrow j}$ have been taken into account ParticleDataGroup2020 to give us:
$n_{j}\left(\mu_{j},T\right)={n_{j}\left(\mu_{j},T\right)}^{prim}+\sum_{k\neq j}{\Gamma_{k\rightarrow j}n_{k}\left(\mu_{k},T\right)}$ (31)
Where $\Gamma_{k\rightarrow j}$ is the branching ratio for the decay of the $k^{th}$ resonance into the $j^{th}$ particle state. The superscript $prim$ stands for the primary number density without decay contribution. The summation is over all the heavier resonances $k$ that decay to the $j$th hadron. Here we have considered the decay contributions to the final state hadron multiplicities from single weak decays of the heavy hadronic resonances as well as those double decays where a weak decay is followed by a strong decay.
The parameters $a$ and $b$ of the model can be adjusted to reproduce the ground state properties of cold nuclear matter (i.e., $T\approx 0\,\text{MeV}$) Vovchenko2015V ; Bethe1971 ; Cleymans1986Q ; Wang2022 ; Greiner1965 ; Bugaev2000 ; Zhang1992 by using the binding energy $E_{B}\approx 16\,\text{MeV}$ and the nucleon number density $n=n_{0}\approx 0.16\,\text{fm}^{-3}$. Following this, one finds $a=329\,\text{MeV}\,\text{fm}^{3}$ and $b=3.42\,\text{fm}^{3}$, which gives the nucleonic hard-core radius $r=0.59\,\text{fm}$, a very reasonable value Vovchenko2016V .
III RESULTS AND DISCUSSION
In this section, we present the outcome of our analysis of particle production, including the strangeness imbalance effect observed in URNNC over a wide range of collision energy, using the above discussed VDW-HRG model. For the $i^{\text{th}}$ hadron having baryon number $B_{i}$, strangeness $S_{i}$, and electric charge $Q_{i}$, we associate baryon chemical potential $\mu_{B}$, strange chemical potential $\mu_{S}$, and electric chemical potential $\mu_{Q}$, respectively. These are treated as model parameters which control the net baryon, strangeness, and electric charge content of the system. The overall chemical potential of the $i^{\text{th}}$ hadron is thus defined as:
$\mu_{i}=B_{i}\mu_{B}+S_{i}\mu_{S}+Q_{i}\mu_{Q}$ (32)
The strangeness conservation criteria and the initial charge-to-baryon ratio of the colliding system can be used to fix $\mu_{S}$ and $\mu_{Q}$, respectively, as these quantities remain conserved during the formation and subsequent evolution of the hot hadronic system created in URNNC Biswas2020 . We can thus write the following Chatterjee2015 :
$\sum_{i}n_{i}\left(T,\mu_{B},\mu_{S},\mu_{Q}\right)S_{i}=\text{Net }S=0$ (33)
$\sum_{i}n_{i}\left(T,\mu_{B},\mu_{S},\mu_{Q}\right)Q_{i}=\mathcal{R}\sum_{i}n_{i}\left(T,\mu_{B},\mu_{S},\mu_{Q}\right)B_{i}$ (34)
Here, =Net $\mathcal{R}=\text{Net }Q/\text{Net }B$ is the net charge-to-net baryon number ratio of the colliding nuclei. For nucleus-nucleus collisions like Au+Au, Pb+Pb, Ag+Ag, etc., we find $\mathcal{R}=N_{p}/(N_{p}+N_{n})\approx 0.4$, with $N_{p}$ and $N_{n}$ denoting the number of protons and neutrons in the colliding nuclei. The NetS is total strangeness of the system.
The freeze-out parameters of the system formed in URNNC, such as temperature and baryon chemical potential, are influenced by various factors, especially the collision energy Tiwari2013 hence, to describe the variations in particle ratios across a wide range of energies up to the LHC, we need to understand how the baryon chemical potential and temperature at the thermo-chemical freeze-out may depend on the collision energies. We have adopted a standard methodology used to fix the temperature $T$ and the baryon chemical potential $\mu_{B}$ of the system for different center-of-mass energies, $\sqrt{s_{NN}}$. Numerous studies Bashir2015 ; Cleymans2006C ; Mishra2008 ; Tiwari2013 ; Andronic2006 ; Cleymans1999 ; Chatterjee2015 ; Biswas2020 ; Tawfik2012P have suggested a suitable ansatz-like dependence of $T$ and $\mu_{B}$ on $\sqrt{s_{NN}}$. Therefore, following the same pattern, we employ the following parameterization Cleymans2006C :
$\mu_{B}=\frac{c}{1+d\sqrt{s_{NN}}}$ (35)
$T=e-f\mu_{B}^{2}-g\mu_{B}^{4}$ (36)
As protons (antiprotons) are the lightest and most abundantly produced baryons (antibaryons) in the system hence we have first attempted to determine the parameters in Eq. 35 and 36 from the best fit to the experimental data on the $\bar{p}/p$ ratio as a function of $\sqrt{s_{NN}}$ in URNNC . The freeze-out temperature has been found to be almost universal at RHIC and LHC, thereby suggesting that chemical freeze-out might occur soon after hadronization Bashir2015 . These studies highlight the complex interplay of factors in determining freeze-out parameters, hence requiring detailed analysis. To ascertain the values of the parameters ($c,d,e,f,g$), we use the best fit criteria, i.e., by obtaining the minimum value of the chi-squared per degree of freedom ($\chi^{2}/\text{dof}$), to the experimentally measured $\bar{p}/p$ ratio distribution over a wide range of collision energies, $\sqrt{s_{NN}}$. The chi-square ($\chi^{2}/\text{dof}$) is computed as:
$\frac{\chi^{2}}{\text{dof}}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left(R_{i}^{\text{the}}-R_{i}^{\text{exp}}\right)^{2}}{\sigma_{i}^{2}}$ (37)
where $R_{i}^{\text{the}}$ and $R_{i}^{\text{exp}}$ represent theoretical and experimental values, respectively, with $\sigma_{i}$ as the experimental uncertainties in the measured quantities at corresponding $\sqrt{s_{NN}}$, with $N$ as the number of degrees of freedom in the data set. The values of the constants turn out to be:
$c=1402\,\text{MeV},\quad d=0.26\,\text{GeV}^{-1},\quad e=172\,\text{MeV},$
$f=0.34\,\text{MeV}^{-1},\quad\text{and}\quad g=0.015\,\text{MeV}^{-3}$
resulting in a satisfactory fit with a value of $\chi^{2}/\text{dof}=1.67$ with N = 15. The $\sqrt{s_{NN}}$ dependent theoretical curve for the $\bar{p}/p$ obtained using the VDW-HRG model calculations is shown in Fig. 1 . Experimental data has been taken from various available sources Adamczyk2017B ; Ahle1999P ; Adams2002S ; Aggarwal2011S ; Aamodt2010M .
Figure 1: $\bar{p}/p$ dependence on $\sqrt{s_{NN}}$
III.1 LIKE MASS ANTIHYPERON TO HYPERON RATIOS
Experimental data show that all antibaryon-to-baryon ratios increase with $\sqrt{s_{NN}}$ and approach unity as we move towards the LHC energies, where the temperature $(T)$ almost saturates at 170 MeV and the chemical potential $(\mu_{B}$) nearly vanishes, indicating an almost baryon-antibaryon symmetric matter. At lower energies (i.e., SPS and AGS), the system maintains a large excess of baryons over antibaryons, resulting in all antibaryon-to-baryon ratios being significantly lower than unity Grebieszkow2004 . In Figs. 4-4, we have shown the energy dependence of antihyperon to hyperon ratios, i.e., $\frac{\bar{\mathrm{\Lambda}}}{\Lambda}$, $\frac{\bar{\mathrm{\Xi}}}{\mathrm{\Xi}}$, and $\frac{\bar{\mathrm{\Omega}}}{\mathrm{\Omega}}$. The minimum $\chi^{2}/\text{dof}$ values for the above three cases are 6.19, 0.81, and 3.43, respectively. The large value for the $\frac{\bar{\mathrm{\Lambda}}}{\Lambda}$ case is apparently due to very small error bars for the lower energy data points.
Figure 2:
Figure 3:
Figure 4:
Figs. 4-4 Variation of Λ¯/$\bar{\Lambda}/\Lambda$, Ξ¯/$\bar{\Xi}/\Xi$ and Ω¯/$\bar{\Omega}/\Omega$ with $\sqrt{s_{NN}}$
The increasing antibaryon to baryon ratios with $\sqrt{s_{NN}}$ also suggests a higher degree of nuclear transparency in the URNNC at higher collision energies, where the bulk of the secondary matter is formed between the two receding nuclei due to high excitation of vacuum and hence is almost baryon symmetric. This effect becomes more prominent as we move towards the highest RHIC and further up to the LHC energies. In lower energy collisions, the bulk secondary matter contains a large contribution from the participant nucleons due to significant baryon stopping in these collisions. This observation is also supported by the decreasing value of net-protons ($p-\bar{p}$) measured in the experiments with increasing collision energy. The matter formed at lower RHIC energies indicates a situation where partial transparency exists in the nuclear collisions. This observation can provide insights into the evolving collision dynamics with increasing energy Abelev2010 . The curves obtained by our VDW-HRG model calculations exhibit good agreement with experimental data Afanasiev2002 ; Antinori2004 ; Appelshauser1998 ; Alt2005Omega ; Mischke2002 ; Aggarwal2011S ; Abbas2013b ; Abelev2010 . The energy-dependent behavior of antihyperon/hyperon ratios, depicted in the Figs. 4, 4, and 4 reveal a mass-hierarchical trend, with a saturation value of 1 being reached earlier for more massive hyperon species. This phenomenon underscores a preference for almost symmetric antihyperons-hyperons production in the case of multistrange hyperons, where the former, i.e., the $\Lambda$, contains only two light valence quarks, while the $\Xi$ contains only one light valence quark, the likely remnants from the initial nucleons. In the case of $\mathrm{\Omega}$, all (strange) valence quarks are newly generated Andronic2006 ; Bashir2015 . The generally unsatisfactory quality of thermal fits to the multi-strange baryons data at SPS energies cannot solely be attributed to these cases, as we shall see below Andronic2006 . These ratios (including $K^{-}/K^{+}$) increase sharply with respect to $\sqrt{s_{NN}}$ and then almost saturate at higher energies (above 200 GeV), reaching a value equal to 1.0 at LHC energy. This behavior shows that the production rates of anti-particles relative to particles continuously increase with increasing $\sqrt{s_{NN}}$ and become almost equal at LHC energy. The excellent agreement between our VDW-HRG model results and the experimental data validates the ability of the model to explain these particle ratios.
III.2 MESONIC RATIOS
Extending our analysis further to obtain the minimum $\chi^{2}/\text{dof}$ fit to the $K^{-}/K^{+}$ ratio data against $\sqrt{s_{NN}}$, we find that it can be best fitted only with a modified (lesser) value of the chemical freeze-out temperature where the parameter $e$ turns out to be 162 MeV instead of 172 MeV (obtained for baryons) while the remaining parameters remain unchanged. This seems to arise due to the later freeze-out of Kaons (including pions as we will see later) than baryons when the system has further cooled down due to expansion. The baryons therefore appear to freeze-out slightly earlier than the kaons due to their larger mean free path in the system. The best fit to the $K^{-}/K^{+}$ data gives $\chi^{2}/\text{dof}=1.28$. This clearly indicates the existence of two chemical freeze-out stages in the system, where the second freeze-out corresponds to the situation where mesons ($K^{-}$, $K^{+}$ and $\pi$) decouple from the system. In some other analyses, it has been suggested that using two chemical freeze-out temperatures improves the thermal model accuracy thereby enhancing its predictive power and better matching with experimental data Assis2013 ; Duarte2012 . Thus, the freeze-out temperature is an effective thermodynamic quantity linked to the system’s expansion.
The $\sqrt{s_{NN}}$ dependent theoretical curve for the $K^{-}/K^{+}$ ratio obtained using the VDW-HRG model calculations is shown in Fig 5 which is in good agreement with the experimental data Abelev2010 ; Abelev2009 ; Acharya2020 ; Adams2002S ; Abelev2013 ; Mischke2002 . The decline in the $K^{-}/K^{+}$ ratio as collision energy decreases is attributed to a rise in net baryon density, favoring the production of $K^{+}$ (having $u$ and $\bar{s}$ valence quarks) over $K^{-}$ (having $\bar{u}$ and $s$ valence quarks).
Figure 5: $\sqrt{s_{NN}}$ dependence of $K^{-}/K^{+}$
As the collision energy increases, the $K^{-}/K^{+}$ ratio exhibits a systematic rise, indicative of the pair production processes for $K^{+}$ and $K^{-}$ at higher energies in almost equal proportion through the reaction channels involving baryons and antibaryons in an almost baryon symmetric matter (having $\mu_{B}\sim 0$). The associated production mechanism yields $K^{+}$ through reactions like $\pi+N\rightarrow Y+K^{+}$ and the production of $K^{-}$ takes place through $\pi+\bar{N}\rightarrow\bar{Y}+K^{-}$, where $N$ represents a nucleon and $Y$ denotes a singly strange hyperon. On the other hand, reactions like $N+\bar{N}\rightarrow K^{-}+K^{+}$ and $\pi^{+}+\pi^{-}\rightarrow K^{-}+K^{+}$ will produce $K^{+}$ and $K^{-}$ in equal proportion Das2014 .
Despite the success of the HRG Model in describing many particle ratios, some unresolved issues still remain while describing the yields of the strange (antistrange) particles relative to the non-strange ones. For nearly a decade, the $K/\pi$, $\Lambda/\pi^{-}$, etc., ratios were not accurately described Beutler2008 ; Tawfik2016 . Pions being the lightest hadrons are produced copiously and carry the bulk of the entropy of the system. The dependence of the $K/\pi$ on $\sqrt{s_{NN}}$ has always been a subject of interest as it is a crucial indicator of the strangeness content of the system relative to its entropy in heavy-ion collisions and has implications for the quark-hadron phase transition. The observed increase and subsequent saturation of the $K/\pi$ ratio with increasing beam energies align with established trends and theoretical expectations Andronic2006 . Assuming simultaneous freeze-out for kaons and pions in our calculation, we have obtained our best fit to the experimental data on $K^{-}/\pi^{-}$ ratio dependence on $\sqrt{s_{NN}}$. The best-fitted curve obtained is shown in Fig 6 in blue color. It is obvious that the theoretical curve provides an unsatisfactory fit to the experimental dataAdams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 . However, such large disagreement has been reported earlier due to strangeness imbalance Castorina2016 ; Munzinger2003 ; Becattini2005 ; Bugaev2016 .
Figure 6: Variation of $K^{-}/\pi^{-}$ with $\sqrt{s_{NN}}$
The theoretical curve is seen to under-predict the experimental data at lower collision energies ($\sqrt{s_{NN}}<10\,\text{GeV}$) and over-predict it at higher energies. We thus introduce a strangeness imbalance factor $\xi_{s}$ which is defined at different energies (say for the $i^{th}$) as followsTawfik2021H ; Biswas2020 :
$\xi_{s}^{i}=\frac{(K^{-}/\pi^{-})_{\text{exp}}^{i}}{(K^{-}/\pi^{-})_{\text{theo}}^{i}}$ (38)
Further, in order to make it convenient to apply this factor to other particle ratios as well for any given collision energy, we have used a parameterization method to represent the energy dependence of $\xi_{s}$. We have used the following ansatz to best represent the obtained $\sqrt{s_{NN}}$ dependent values of $\xi_{s}$:
$\xi_{s}=\alpha+\beta\left(1+\left[\frac{\sqrt{s_{NN}}}{\gamma}\right]^{\delta}\right)^{-1}$ (39)
The best spline fit provides $\alpha=0.49$, $\beta=8.37\times 10^{6}$, $\gamma=0.26\,\text{GeV}$, and $\delta=4.63$.
Figure 7: $\sqrt{s_{NN}}$ dependence of $\xi_{s}$
In Fig 7, we have shown the energy dependence of $\xi_{s}$. Utilizing the ansatz of Eq. 39, we have redefined the strangeness imbalance corrected strange (antistrange) particle number densities in our calculation at different collision energies as
$n_{s}^{i}\left(\text{corrected}\right)=\left[\xi_{s}^{i}\right]^{h}n_{s}^{i}\left(\text{uncorrected}\right)$ (40)
The quantity $h$ represents the strangeness content of a given particle. For singly strange particles (i.e., kaons, antikaons, lambda, antilambda, sigma, antisigma) we have $h=1$; for doubly strange particles (i.e., cascades, anticascades) we have $h=2$; and similarly for triply strange particles (i.e., omega, antiomega) we have $h=3$. In Fig 6, the corrected theoretical curve for the $K^{-}/\pi^{-}$ ratio using the above ansatz in Eq. 39 is shown by the red curve, which has a very small $\chi^{2}/\text{dof}=0.04$, resulting from the correction procedure. Moreover, using the values of $\xi_{s}$ obtained from Eq. 39, we have also calculated the strangeness imbalance corrected $K^{+}/\pi^{+}$ ratio for different collision energies. In Fig 8, the theoretical curve for the $K^{+}/\pi^{+}$ ratio is shown in red, giving a reasonable value of $\chi^{2}/\text{dof}=0.16$. Interestingly, the theoretical curve also explains the horn structure at lower energies, $\sqrt{s_{NN}}\sim 7-8\,\text{GeV}$. The steep variation of the $(K^{+}/\pi^{+})$ and $(K^{-}/\pi^{-})$ ratios at lower energies reflects the threshold for strangeness production affected by the rapid initial increase in the temperature with $\sqrt{s_{NN}}$. A steep drop in the $K^{+}/\pi^{+}$ near the horn may arise due to the rapidly falling baryon chemical potential, as baryon-rich matter favors $K^{+}$ production, as discussed above Adams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 ; RecentRF ; NA61SHINE2020 ; Ahle1999P .
Figure 8: Variation of $K^{+}/\pi^{+}$ with $\sqrt{s_{NN}}$
Figure 9: Dependence of $\pi^{-}/\pi^{+}$ and $K^{-}/K^{+}$ on $\sqrt{s_{NN}}$
In Fig. 9, we have shown the energy dependence of the $\pi^{-}/\pi^{+}$ ratio. The green curve shows the theoretical fit, which matches the experimental data quite well with $\chi^{2}/\text{dof}=0.37$. It may be recalled here that in the calculation of all particle ratios, we have also applied the criteria of conserved charge to baryon ratio in the experiments in order to fix the electric chemical potential $\mu_{Q}$. We find that the $\pi^{-}/\pi^{+}$ ratio decreases with $\sqrt{s_{NN}}$ gradually and saturates towards unity as the system tends to become almost charge symmetric. In the same figure, we have again shown the $K^{-}/K^{+}$ ratio for comparison, as both ratios become equal ($\sim 1$) for very large $\sqrt{s_{NN}}$. At low beam energies, the $\pi^{-}/\pi^{+}$ ratios exhibit values larger than unity, possibly due to significant contributions to the positive charges from protons and other resonances, particularly the $\Delta$ baryons. For the ease of comparison we have also shown the $K^{-}/K^{+}$ ratio in the same figure. The distribution characteristics of the $\pi^{-}/\pi^{+}$ and $K^{-}/K^{+}$ ratios differ from each other however both ratios are seen approaching 1 due to nuclear transparency effects. The theoretical results are in reasonable agreement with the experimental data Bratkovskaya2004 ; Adams2002S ; Abelev2010 ; Abelev2009 ; Abelev2013 ; Adamczyk2017B ; Adler2002 . This also explains the observation that the $K^{-}/\pi^{-}$ and $K^{+}/\pi^{+}$ ratios in Figs. 6 and 8 approach the same value ($\sim 0.15$) towards the highest RHIC and LHC energies.
In Figs. 10, 11, 12, 13, and 14, we have shown the results of our calculations along with the experimental data for several particle ratios, i.e., $\frac{\Lambda(\bar{\Lambda})}{\pi^{-}}$, $\frac{\Xi(\bar{\Xi})}{\pi^{-}(\pi^{+})}$, $\Omega/\pi$, $\Phi/\pi^{-}$, and $\Phi/K^{-}$. We find that a reasonable fit is obtained in all these cases, thus highlighting the need for consideration of the strangeness imbalance effect in the system. Corresponding experimental data are taken from the references Adam2020 ; Abelev2010 ; Adamczyk2017B ; Antii2003 ; RecentRF ; Alt2008 ; Aggarwal2011S ; Friese2004 ; E9172003 . The $\mathrm{\Phi}\left(s\bar{s}\right)$ mesons with hidden strangeness, having relatively small cross sections with other hadrons, are assumed to freeze-out earlier than Kaons and pions (i.e., along with baryons). Our model calculations are able to describe the $\phi/\pi$ and $\Omega/\pi$ ratios reasonably well. Our results are in better agreement than an earlier result Tiwari2013 . Due to the strangeness imbalance effect the like-mass anti-hyperon to hyperon as well as antikaon to kaon ratios are not altered, as their abundances in the system are equally affected. In their theoretical calculations the correction factor $\xi_{s}$ cancels.
Figure 10:
Figure 11:
Figure 12:
Figure 13:
Figure 14:
Figs. 10-14 Variation of $\Lambda(\bar{\Lambda})/\pi^{-}$, $\Xi/\pi$ , $\Omega/\pi^{-}$, $\phi/\pi^{-}$ and $\phi/K^{-}$ with $\sqrt{s_{NN}}$
As the protons (antiprotons) are the lightest baryons (antibaryons) hence most high-mass baryons (antibaryons) decay into protons (antiprotons). The $p/\pi$ ratio, therefore, characterizes well the overall baryon (antibaryon) production relative to the total particle multiplicity. To understand this, we have shown the $p/\pi^{+}$ and $\bar{p}/\pi^{-}$ ratios energy dependence in Fig 15. The theoretical curve appears to fit the data quite well with $\chi^{2}/\text{dof}=1.60$ and $4.20$, respectively. The larger value for the second case arises due to very small error bars in the experimental data Nandi2020 ; Aggarwal2011S ; Adamczyk2017B ; Abelev2010 .
Figure 15: Variation of $p/\pi$ with $\sqrt{s_{NN}}$
Figure 16: Correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios
We have analyzed the correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios in URNNC over a wide energy range from 5 GeV to 2.76 TeV. The correlation is best described by a power-law type relationship as $K^{-}/K^{+}=\left(\bar{p}/p\right)^{\nu}$. This can serve as validation for any thermal model as it indicates how kaon production relates to net baryon density. In Fig 16, we have shown the experimental data points along with curves obtained from different approaches. The blue curve shows the result of our calculations. The pink curve is obtained by considering the light valence quark compositions, thus providing $\nu\approx 0.33$ but fits the data somewhat poorly. For comparison, if we represent the theoretical blue curve by a power-law fit, this provides $\nu\approx 0.18$, which contrasts sharply with the light valence quark composition approach. However, if we obtain a direct fit to the data using the power law, we get $\nu\approx 0.21$, which is quite close to our theoretical results. This correlation is therefore well described by the VDW-HRG type approach in the framework of the grand canonical ensemble-based thermal model. Experimental data at different energies are taken from the references Bearden2003 ; Stiles2006 ; Das2014 ; Abelev2009 .
III.3 ENERGY DEPENDENCE OF THERMAL FREEZE-OUT PARAMETERS ($T$, $\mu_{B}$, $\mu_{s}$)
In order to highlight the energy dependence of the freeze-out parameters of the model, we have shown in Figs. 17 and 18 the $\sqrt{s_{NN}}$ dependence of $T$, $\mu_{B}$, $\mu_{s}$, and $\mu_{Q}$. The energy dependence of $\xi_{s}$ has already been discussed in the previous section shown in Fig 7. Fig. 17 shows how the system’s temperature varies with collision energy in URNNC. The two separate curves are for baryonic and mesonic degrees of freedom as they are found to freeze-out at different stages of the system’s evolution as discussed earlier also. However, both curves saturate beyond $\sqrt{s_{NN}}\sim 60\,\text{GeV}$ at 172 MeV and 162 MeV, respectively. This may indicate the onset of nuclear transparency around this energy, with lesser stopping of the colliding nuclear medium and a large production of secondary hadrons during the evolution of the system Bashir2015 . In Fig. 18, we have shown the energy dependence of $\mu_{B}$ and $\mu_{s}$ extracted from the thermal fits of particle ratios. The $\mu_{B}$ is seen to initially decrease rapidly with $\sqrt{s_{NN}}$ and later approach a negligibly small value beyond 200 GeV. The $\mu_{s}$, which is obtained by applying the strict strangeness conservation criteria in the hot and dense system formed in URNNC, initially shows a slight increasing trend up to 10 GeV, reaching a value of approximately 50 MeV, and then shows a decreasing trend to approach almost zero value beyond 200 GeV. Both $\mu_{B}$ and $\mu_{s}$ thus approach zero as the matter becomes more and more baryon symmetric with increasing $\sqrt{s_{NN}}$, which again is an indicator of the nuclear transparency effect setting in at the higher RHIC energies and beyond up to LHC energies.
Fig. 19 illustrates a smooth parabolic dependence of the chemical freeze-out stages in the $T-\mu_{B}$ plane. The vertical axis represents the chemical freeze-out temperature, while the horizontal axis shows the baryon chemical potential, both obtained using our model calculations for different $\sqrt{s_{NN}}$. The red and blue curves in the figure again correspond to the baryonic and mesonic degrees of freedom, respectively. At RHIC energies, the chemical freeze-out temperatures $T_{che}\left(\approx 163-170\,\text{MeV}\right)$ are close to the hadron-QGP phase transition temperature predicted by lattice gauge theory Bashir2015 ; Baban2013 ; Tiwari2013 thus indicating that freeze-out occurs almost immediately after the phase transition.
Our model results align closely with the findings from some previous studies on freeze-out parameters in heavy-ion collisions Cleymans2006C ; Munzinger2001 ; Adams2005E ; Baran2003 ; Florkowski2001fp ; Chatterjee2015 ; Alba2020ChemicalFP . The freeze-out temperatures ($T$) and baryon chemical potentials ($\mu_{B}$) for RHIC energies have been earlier reported, showing a range of $T$ values from 163 MeV to 177 MeV and $\mu_{B}$ values from 24 MeV to 29 MeV for $\sqrt{s_{NN}}=200\,\text{A GeV}$, and $T$ values from 165 MeV to 174 MeV and $\mu_{B}$ values from 38.1 MeV to 46 MeV for $\sqrt{s_{NN}}=130\,\text{A GeV}$ Munzinger2001 ; Cleymans2005C .
Figure 17:
Figure 18:
Figs. 17-18 Variation of $T$ and $\mu_{B}$, $\mu_{s}$ with $\sqrt{s_{NN}}$
Figure 19: Variation in the $T-\mu_{B}$ plane
IV SUMMARY AND CONCLUSION
In summary we have systematically applied the statistical VDW-HRG model in the framework of grand canonical ensemble to analyze relative particle yields in ultra-relativistic nucleus-nucleus collisions. We have used the values of the model parameters representing the attractive and repulsive parts of interaction, which are obtained from the ground state properties of nuclear matter. Using an established ansatz form for the chemical freeze-out temperatures and baryon chemical potential, we have determined the thermo-chemical freeze-out parameters from the $\bar{p}/p$ and $K^{-}/K^{+}$ ratios. We have found that there is a need to consider separate freeze-out temperatures for baryons and mesons (Kaons and pions). This is justified by their different interaction cross-sections in the system. We have further explained strange particle-to-pion ratios successfully, including the experimentally observed hump in $K^{+}/\pi^{+}$, by introducing a strangeness imbalance factor ($\xi_{s}$) derived from $K^{-}/\pi^{-}$ ratios. Using the model, we have quite successfully explained the experimental results for several like and unlike mass hadron ratios over a wide range of collision energies. The experimentally observed correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios is well explained by our model calculations. A common feature of all ratios is that they tend to saturate at large collision energies. This is a clear indication of a nearly constant temperature and chemical potential (though vanishingly small) at higher collision energies, i.e., beyond 200 GeV. In the case of unlike mass particle ratios, the initially increasing temperature with $\sqrt{s_{NN}}$ favors a relatively rapid increase in the production of heavier particles. The decreasing chemical potential also favors the production of antibaryons, while for baryons, it does the opposite. We have illustrated the energy dependence of thermal freeze-out parameters $T$, $\mu_{B}$, and $\xi_{s}$, indicating the complex nature of particle production at high energies.
ACKNOWLEDGMENTS
We would like to thank Jajati K. Nayak and Abdel Nasser Tawfik for their help in providing well-structured experimental data and their suggestions from time to time. The authors are also thankful to Jan-E Alam, Sandeep Chatterjee and Sabyasachi Ghosh for their valuable suggestions.